OpenStudy (anonymous):

lim of x / ( sqrt( x + 1 ) - 1 = 2 x -> 0 x approaching to 0, the limit is 2 how

OpenStudy (anonymous):

\[\LARGE \lim_{x\to 0}\frac{x}{\sqrt{x+1}-1}=2\]

OpenStudy (anonymous):

yah. why lim = 2 ? how

OpenStudy (anonymous):

I'm confused too O_O... :(

OpenStudy (anonymous):

omg... :D it's alright :D

OpenStudy (experimentx):

put the value of x=0

OpenStudy (anonymous):

experiment, then we don't get 2 never

OpenStudy (experimentx):

@Kreshnik how do you make that appear large??

OpenStudy (experimentx):

try and see what you get??

OpenStudy (anonymous):

then we get 0

OpenStudy (experimentx):

yes that is the answer.

OpenStudy (anonymous):

no the limit is = 2 . i have answer already. i have a graphic application that also indicates limit to be 2

OpenStudy (shayaan_mustafa):

i am getting answer to your question 1/2 not 0

OpenStudy (experimentx):

then ... i must see ,,, OO ... didn't realized the indeterminate form.

OpenStudy (experimentx):

multiply both by conjugate and see what happens/

OpenStudy (anonymous):

but i want the answer to be 2. :) :)

OpenStudy (anonymous):

@experimentX \[\huge \lim_{x\to 0} \frac{\text{Use \huge }}{\text{=this}}\] lim_{x \to 0} \frac{x}{\sqrt{x+1}-1}=2 \[\lim_{x \to 0} \frac{x}{\sqrt{x+1}-1}=2\] \large lim_{x \to 0} \frac{x}{\sqrt{x+1}-1}=2 \[\large \lim_{x \to 0} \frac{x}{\sqrt{x+1}-1}=2\] \Large lim_{x \to 0} \frac{x}{\sqrt{x+1}-1}=2 \[\Large \lim_{x \to 0} \frac{x}{\sqrt{x+1}-1}=2\] \LARGE \lim_{x \to 0} \frac{x}{\sqrt{x+1}-1}=2 \[\LARGE \lim_{x \to 0} \frac{x}{\sqrt{x+1}-1}=2\] \huge \lim_{x \to 0} \frac{x}{\sqrt{x+1}-1}=2 \[\huge \lim_{x \to 0} \frac{x}{\sqrt{x+1}-1}=2\] @experimentX

OpenStudy (experimentx):

yes ... multiplying by conjugate will give limit 2

OpenStudy (anonymous):

yah experiment u r right ...

OpenStudy (shayaan_mustafa):

but applying L'hopitals give 1/2 :s

OpenStudy (experimentx):

using L'hospital rule will also give the same 2 ... you might be forgetting something.

OpenStudy (anonymous):

@experimentX please tell me how do you use L'H'R here.. both sides? :O :$

OpenStudy (experimentx):

differentiate both function ... on top and bottom

OpenStudy (anonymous):

experiment, how did u know that we have to multiply by denominator's conjugate?

OpenStudy (experimentx):

experience man ... thats one of standard tricks.

OpenStudy (anonymous):

\[ \lim_{x \rightarrow 0} \sin x / x\]

OpenStudy (anonymous):

i need it's limit too

OpenStudy (experimentx):

\[ \huge \lim_{x \to 0} \frac{\frac{d(x)}{dx}}{\frac{d(\sqrt{x+1}-1)}{dx}}=2 \] \[ \huge \lim_{x \to 0} \frac{1}{\frac{1}{2\sqrt{x+1}}-0}=2 \]

OpenStudy (anonymous):

hahaaaahh .. NOW we're talking ;) well done !

OpenStudy (experimentx):

@Kreshnik put the value of x=0 on the top ... that is LHR @jatinbansalhot it's value is 1 .... it's quite complicated => one is geometrical proof other proof is quite complicated.

OpenStudy (anonymous):

i have a new question plz try that one i urgently need that

OpenStudy (experimentx):

@jatinbansalhot the best thing you can do to view it's proof is view MIT video lectures.

OpenStudy (anonymous):

i had posted that

OpenStudy (anonymous):

Experiment which lectures single variable calculus

OpenStudy (anonymous):

single variable calculus is calculus 1 ?

OpenStudy (anonymous):

send me exact link of those lectures plz .

OpenStudy (experimentx):

i am beginner ... there's now way i could do that. @jatinbansalhot yes ... on single variable calculus. it's beautifully done there ... geometrically.

OpenStudy (anonymous):

i'm doing Calculus 1 . Is single variable calculus same ?

OpenStudy (experimentx):

yes ... lecture 3

OpenStudy (anonymous):

only lecture 3 ? plz send me the link